[clas_members] RC daily summary for Sunday February 26

silvia at jlab.org silvia at jlab.org
Sun Feb 26 10:25:39 EST 2023

Day shift (Saturday):
Taken Moeller run, polarization a bit lower than usual: 77.972 +/- 1.464 %.
Reconfigured things for production running, and started taking data on ND3
at 4nA, while the target was displaying a record polarization of -51%.
One FTOF channel tripped and could not be recovered, expert informed, the
channel has been turned off for now (S6 P1a 23 R).
71M events collected.

Swing shift (Saturday):
Continued data taking on ND3 at 4nA.
Some HV alarms for ECAL and some BMT trips, but otherwise a smooth shift,
114M events collected.
The target polarization has been decreasing since it started getting beam,
it is around -36% at the end of the shift. The target group is watching it
and they periodically intervene tuning the MW to try to slow down the
decay (thank you!).

Owl shift (Sunday):
Continuing data taking on ND3 at 4 nA.
Some minor problems with DAQ and drifting beam positions, but overall a
smooth shift with 108M events collected.

Day shift:
Smooth data taking so far. We will switch to HWP OUT at around 2PM this

Tentative plan:
As the scheduled beam downtime will be next Wednesday, and
anneal+repolarization takes a long time, it was proposed on Friday at the
RC meeting to change to a background target (C and/or empty), which is
fast, run for a day or so this way, and then use Wednesday to put back ND3
and repolarize it positive. This would be the last ND3 data taking for
this experiment, and then we can switch permanently to NH3 for the
remained of RGC. This change could happen either Monday evening or Tuesday
morning, depending on how well the target holds (right now it is around
Is this OK with everybody? We can discuss by email, I don't think we need
a meeting for today.

More information about the clas_members mailing list